# Near Coincidences and Nilpotent Division Fields

### v1

@@ 5. Nilpotent groups of prime-power level
Suppose that $p$ is a prime and that $E/\mathbb{Q}$ is an elliptic curve such that $\mathbb{Q}(E[p^{k}])/\mathbb{Q}$ is a nilpotent extension for some $k\geq 2$. The first observation we make is that since nilpotency is closed under quotients, Proposition 4.4, we know that $\mathbb{Q}(E[p^{i}])/\mathbb{Q}$ is a nilpotent extension for all $1\leq i\leq k,$ in particular, we this would mean $\mathbb{Q}(E[p])/\mathbb{Q}$ is a nilpotent extension. As usual, we will have to handle the case when $p=2$ separately, but thanks to Proposition 4.9, when $p$ is odd, we only have to deal with the case that $\operatorname{Im}\bar{\rho}_{E,p}$ is contained in either the normalizer of a split or non-split Cartan subgroup of $\operatorname{GL}_{2}(\mathbb{Z}/p\mathbb{Z}).$
@@ 5.1. The case when $p=2$
The next step is to search for admissible nilpotent subgroups of $G_{4}$ up to conjugation with the additional property that their image mod $2$ is exactly equal to $G_{2}$. A quick search shows that there are no such groups. Thus in the case when $\operatorname{Im}\bar{\rho}_{E,2}$ is conjugate to $G_{2}$, there is no way that $\mathbb{Q}(E[4])/\mathbb{Q}$ can be a nilpotent extension.
@@
and in particular $\pi_{2}^{-1}(G_{2})$ is a 2-group. From Theorem 4.3, we know that $\pi_{2}^{-1}(G_{2})$ is always nilpotent. The upshot of this is that if $E/\mathbb{Q}$ is an elliptic curve a point of order two defined over $\mathbb{Q}$, then for every $k\geq 1$, $\mathbb{Q}(E[2^{k}])/\mathbb{Q}$ is a nilpotent extension.
@@ 5.2. The case when $p$ is odd
Suppose that $G$ is a nilpotent subgroup of $\operatorname{GL}_{2}(\mathbb{Z}/p^{2}\mathbb{Z})$ and let $\pi:G\to\operatorname{GL}_{2}(\mathbb{F}_{p})$ be the reduction mod $p$ map. Assume that $p\nmid|\pi(G)|$. Then at least one of the following is true:
@@
Let $P$ be a Sylow $p$-subgroup of $G$. Since $|\pi(G)|$ has order coprime to $p$, we have that $\pi(P)=\{1\}$. In particular, $P$ is contained in the set of matrices $\equiv I\pmod{p}$. The set of matrices $\equiv I\pmod{p}$ is an abelian subgroup of $\operatorname{GL}_{2}(\mathbb{Z}/p^{2}\mathbb{Z})$ order $p^{4}$. In particular $P$ is abelian. Let
@@
be a complement of $P$ in $G$.
@@
This implies that there is some $X\in M_{2}(\mathbb{F}_{p})$ so that $I+pX\in P$ and $X$ is not a scalar multiple of the identity. If $Y\in\pi(G)$, then since $\pi:H\to\operatorname{GL}_{2}(\mathbb{Z}/p\mathbb{Z})$ is injective, there is some $\tilde{Y}$ with order coprime to $p$ so that $\pi(\tilde{Y})=Y$. The assumption that $G$ is nilpotent implies that $(I+pX)$ must commute with $\tilde{Y}$, and this implies that $XY=YX$ in $M_{2}(\mathbb{F}_{p})$. The assumption on $X$ implies that $X$ is a cyclic matrix. This is a matrix $X$ whose minimal polynomial and characteristic polynomial are the same.
Corollary 4.4.18 of [15] implies that for every cyclic matrix $X$, its centralizer in $M_{2}(\mathbb{F}_{p})$ is equal to $\mathbb{F}_{p}[X]$, the set of all polynomials in $X$ with coefficients in $\mathbb{F}_{p}$. This is a commutative subring of $M_{2}(\mathbb{F}_{p})$, and this implies that $\pi(G)\subseteq\mathbb{F}_{p}[X]^{\times}$ is abelian. Since $\pi:H\to\operatorname{GL}_{2}(\mathbb{F}_{p})$ is injective, it follows that $H$ is abelian. Since $G\cong P\times H$, it follows that $G$ is abelian.
@@
Assume that $E/\mathbb{Q}$ is an elliptic curve, $p$ is an odd prime, and $\mathbb{Q}(E[p^{2}])/\mathbb{Q}$ is nilpotent. Let $G=\operatorname{Im}\bar{\rho}_{E,p^{2}}$. If we are in case (1) of Proposition 5.2, then $\mathbb{Q}(E[p^{2}])/\mathbb{Q}$ is abelian, which contradicts the main result of [14]. If we are in case (2) of Proposition 5.2 and $\pi:G\to\operatorname{GL}_{2}(\mathbb{F}_{p})$ is the reduction mod $p$ map, then $\ker(\pi)\cap(G\cap\operatorname{SL}_{2}(\mathbb{Z}/p^{2}\mathbb{Z}))=1$ and this implies that we have a near coincidence of level $(p^{2},p)$. By Proposition 3.7, we must have that $p=3$ and $E$ corresponds to a rational point on the curve with RSZB label 9.27.0.1. However, [26] implies that for such an elliptic curve, the mod $9$ image Galois must equal 9.27.0.1, which is not nilpotent. ∎
@@ 6. Nilpotent groups of composite level
Next we consider the case that $E/\mathbb{Q}$ is an elliptic curve without complex multiplication under the assumption of Conjecture 1.6. In this case, Section 4 implies that if $\mathbb{Q}(E[p])/\mathbb{Q}$ is nilpotent then either $p\in\{2,5\}$ or $p$ is a Mersenne prime and $\operatorname{Im}\bar{\rho}_{E,p}$ is contained in the normalizer of the non-split Cartan subgroup of $\operatorname{GL}_{2}(\mathbb{F}_{p})$, which entails that $p\in\{2,3,5,7\}$. All that remains is for us to determine which combinations of the images listed in Table 1 that have rational points can occur simultaneously. All that is necessary to construct the modular curves parameterizing the elliptic curves with two given images simultaneously is to take the fiber product of the two curves. In other words, to compute the modular curve parameterizing elliptic curves whose $\pmod{p}$ image is in $G$ and $\pmod{q}$ is in $H$, we simply need to compute the curve given by $\pi_{G}(x)=\pi_{H}(y)$. The results of this computation are listed in the following table.
@@
Table 2. Fiber products of groups from Table 1.
The curves that are genus 1 or 2 with rank zero can be handeled using standard techniques. The rank 0 genus 3 curve was shown to have no non-cuspidal rational points corresponding to elliptic cuvres without complex multiplication in [25]. This leaves us with with one remaining curve, a genus 19 rank 15 curve.
In theory, this curve could be attacked using the method of Chabauty and Coleman since the genus higher than the rank, but computing on a curve of genus 19 is rather unwieldy. In this case, $G=C_{s}^{+}(5)$ and $H=C_{ns}^{+}(7)$, and Theorem 1.4 of [20] states the following.
@@
[20, Theorem 1.4] Let $E/\mathbb{Q}$ be an elliptic cuvre without complex multiplication. Suppose that there exists a prime $q$ for which $\operatorname{Im}\bar{\rho}_{E,q}$ is contained in a subgroup of $C_{s}^{+}(q)$. Then $\bar{\rho}_{E,p}$ is surjective for all $p>37.$
The proof of Theorem 6.1 uses an adaptation of Mazur’s formal immersion argument in [23, 24]. The idea is to show that if there is a curve with some combination of these images, then the $j$-invariant of this curve must be an integer. Then one can use the specifics of the various $j$-maps to show that this can’t happen. Examining the proof of this theorem, the only reason that the assumption that $p>37$ is there is becuase of the relationship to Serre’s uniformity problem. Thus, the techniques used to prove this theorm also work if this condition were changed to $p>5$. For this reason, there is no elliptic curve $E/\mathbb{Q}$ without complex multiplication for which $\mathbb{Q}(E[35])/\mathbb{Q}$ is nilpotent. In Table 3, we give models for the modular curves from Table 2 that have non-cuspidal rational points.
@@ 21.63.1.1 $y^{2}+y=x^{3}+12$ $f_{21}(x,y)=f_{7}\left(\frac{x^{2}+5x-14}{x^{2}-4x+3y+19}\right)$
Table 3. Models of the modular curves in question. The models for curves of prime level come from [33]. The remaining genus 0 curves can be computed as fiber producs of the curves of prime level. The models for the genus 1 modular curves of composite level are computed as the fiber product of the prime level modular curves that are computed in [33] as well.
Working without the assumption of Conjecture 1.6, we must consider the possibility that there is an elliptic curve $E/\mathbb{Q}$ for which $\operatorname{Im}\bar{\rho}_{E,5}$ is contained in the normalizer of the split Cartan mod $5$ and for which $\operatorname{Im}\bar{\rho}_{E,p}$ is contained in the normalizer of a non-split Cartan modulo $p$ for some Mersenne prime $p$. This is possible for $p=3$, and the elliptic curves for which this occurs are parametrized by the modular curve with label 15.45.1.1. For $p>3$, the extension of Theorem 6.1 just mentioned shows that this is impossible.
In addition, we must consider the possibility that there is an elliptic curve $E/\mathbb{Q}$ for which $\mathrm{im}\,\bar{\rho}_{E,p}$ is contained in the normalizer of a non-split Cartan modulo $p$ for some Mersenne prime $p$, and which also has a rational point of order $2$. If this case were to occur, then $\mathbb{Q}(E[2^{k}p])/\mathbb{Q}$ would be nilpotent for all $k\geq 1$. We now use the main result of [19].
Theorem 6.2.
[19, Theorem 1.1] Let $E/\mathbb{Q}$ be an elliptic curve without complex multiplication with a non-trivial cyclic $\mathbb{Q}$-isogeny. Then, for $p>37$, $\bar{\rho}_{E,p}$ is surjective.
This result immediately rules out the possibility that $p>37$. The only Mersenne primes $\leq 37$ are $3$, $7$ and $31$, and the former two cases are considered in Table 2. It remains to consider the case that $p=31$. Proposition 2.1 of [19] shows that if $E/\mathbb{Q}$ is an elliptic curve with a $\mathbb{Q}$-rational $2$-isogeny and there is a prime $p$ for which the image of $\bar{\rho}_{E,p}$ is contained in the normalizer of a non-split Cartan, than $j(E)\in\mathbb{Z}$. Moreover, Theorem 2.3 of [19] shows that there are only finitely many integral $j$-invariants of elliptic curves $E/\mathbb{Q}$ with a $\mathbb{Q}$-rational $2$-isogeny and explicitly lists them. Using data from the LMFDB about mod $p$ images of Galois shows that no elliptic curve $E/\mathbb{Q}$ with one of these $j$-invariants has mod $31$ image of Galois contained in the normalizer of the non-split Cartan. This concludes the proof of Theorem 1.7.

### v3

@@ 5. Nilpotent groups of prime-power level
Suppose that $p$ is a prime and that $E/\mathbb{Q}$ is an elliptic curve such that $\mathbb{Q}(E[p^{k}])/\mathbb{Q}$ is a nilpotent extension for some $k\geq 2$. The first observation we make is that since nilpotency is closed under quotients, Proposition 4.4, we know that $\mathbb{Q}(E[p^{i}])/\mathbb{Q}$ is a nilpotent extension for all $1\leq i\leq k,$ in particular, and this would mean $\mathbb{Q}(E[p])/\mathbb{Q}$ is a nilpotent extension. As usual, we will have to handle the case when $p=2$ separately, but thanks to Proposition 4.10, when $p$ is odd, we only have to deal with the case that $\operatorname{Im}\bar{\rho}_{E,p}$ is contained in either the normalizer of a split or non-split Cartan subgroup of $\operatorname{GL}_{2}(\mathbb{Z}/p\mathbb{Z}).$
@@ 5.1. The case when $p=2$
The next step is to search for admissible nilpotent subgroups of $G_{4}$ up to conjugation with the additional property that their image mod $2$ is exactly equal to $G_{2}$. Such a subgroup must have order which is a multiple of $6$ (a factor of $3$ coming from the image in $G_{2}$ and a factor of $2$ coming from the determinant being surjective). There are two such groups up to conjugacy: one with order $6$ and one with order $12$, with the former contained in the latter. Neither of these groups are admissible. Computing conjugacy classes of the order $12$ subgroup shows that there are three conjugacy classes of elements of order $2$ and none of these elements fix an element of $(\mathbb{Z}/4\mathbb{Z})^{2}$ of order $4$, which the image of complex conjugation under $\bar{\rho}_{E,4}$ must. (Code for this calculation can be found in the file subsec51.m at [8].) Thus in the case when $\operatorname{Im}\bar{\rho}_{E,2}$ is conjugate to $G_{2}$, there is no way that $\mathbb{Q}(E[4])/\mathbb{Q}$ can be a nilpotent extension.
@@
and in particular $\pi_{2}^{-1}(G_{2})$ is a 2-group. From Theorem 4.3, we know that $\pi_{2}^{-1}(G_{2})$ is always nilpotent. The upshot of this is that if $E/\mathbb{Q}$ is an elliptic curve with a point of order two defined over $\mathbb{Q}$, then for every $k\geq 1$, $\mathbb{Q}(E[2^{k}])/\mathbb{Q}$ is a nilpotent extension.
@@ 5.2. The case when $p$ is odd
Suppose that $G$ is a nilpotent subgroup of $\operatorname{GL}_{2}(\mathbb{Z}/p^{2}\mathbb{Z})$ and let $\pi\colon G\to\operatorname{GL}_{2}(\mathbb{F}_{p})$ be the reduction mod $p$ map. Assume that $p\nmid|\pi(G)|$. Then at least one of the following is true:
@@
Let $P$ be a Sylow $p$-subgroup of $G$. Since $|\pi(G)|$ has order coprime to $p$, we have that $\pi(P)=\{1\}$. In particular, $P$ is contained in the set of matrices $\equiv I\pmod{p}$. The set of matrices $\equiv I\pmod{p}$ is an abelian subgroup of $\operatorname{GL}_{2}(\mathbb{Z}/p^{2}\mathbb{Z})$ order $p^{4}$. In particular $P$ is abelian and $\ker\pi=P$. Let
@@
be a complement of $P$ in $G$. Note that $\pi(G)=\pi(PH)=\pi(P)\pi(H)=\{1\}\cdot\pi(H)=\pi(H)$ and also $H\cap\ker\pi\subseteq H\cap P=\{1\}$. Thus $\pi\colon H\to\pi(G)$ is an isomorphism.
@@
This implies that there is some $X\in M_{2}(\mathbb{F}_{p})$ so that $I+pX\in P$ and $X$ is not a scalar multiple of the identity. If $Y\in\pi(G)$, there is some $\tilde{Y}\in H$ so that $\pi(\tilde{Y})=Y$. The assumption that $G$ is nilpotent implies that $(I+pX)$ must commute with $\tilde{Y}$, and this implies that $XY=YX$ in $M_{2}(\mathbb{F}_{p})$. The assumption on $X$ implies that $X$ is a cyclic matrix. This is a matrix $X$ whose minimal polynomial and characteristic polynomial are the same.
Corollary 4.4.18 of [17] implies that for every cyclic matrix $X$, its centralizer in $M_{2}(\mathbb{F}_{p})$ is equal to $\mathbb{F}_{p}[X]$, the set of all polynomials in $X$ with coefficients in $\mathbb{F}_{p}$. This is a commutative subring of $M_{2}(\mathbb{F}_{p})$, and this implies that $\pi(G)\subseteq\mathbb{F}_{p}[X]^{\times}$ is abelian. Since $H\simeq\pi(G)$, it follows that $H$ is abelian. Since $G\simeq P\times H$, it follows that $G$ is abelian.
@@
Assume that $E/\mathbb{Q}$ is an elliptic curve, $p$ is an odd prime, and $\mathbb{Q}(E[p^{2}])/\mathbb{Q}$ is nilpotent. Let $G=\operatorname{Im}\bar{\rho}_{E,p^{2}}$. If we are in case (1) of Proposition 5.2, then $\mathbb{Q}(E[p^{2}])/\mathbb{Q}$ is abelian, which contradicts the main result of [16]. If we are in case (2) of Proposition 5.2 and $\pi\colon G\to\operatorname{GL}_{2}(\mathbb{F}_{p})$ is the reduction mod $p$ map, then $\ker(\pi)\cap(G\cap\operatorname{SL}_{2}(\mathbb{Z}/p^{2}\mathbb{Z}))=1$ and this implies that we have a near coincidence of level $(p^{2},p)$. By Proposition 3.11, we must have that $p=3$ and $E$ corresponds to a rational point on the curve with RSZB label 9.27.0.1. However, the main result of [33] implies that for such an elliptic curve, the mod $9$ image of Galois must equal 9.27.0.1, which is not nilpotent. ∎
@@ 6. Nilpotent groups of composite level
Next we consider the case that $E/\mathbb{Q}$ is an elliptic curve without complex multiplication under the assumption of Conjecture 1.8. In this case, Section 4 implies that if $\mathbb{Q}(E[p])/\mathbb{Q}$ is nilpotent then either $p\in\{2,5\}$ or $p$ is a Mersenne prime and $\operatorname{Im}\bar{\rho}_{E,p}$ is contained in the normalizer of the non-split Cartan subgroup of $\operatorname{GL}_{2}(\mathbb{F}_{p})$, which entails that $p\in\{2,3,5,7\}$. All that remains is for us to determine which combinations of the possible nilpotent mod $p$ images can occur simultaneously. Given possible mod $p$ and mod $q$ images of Galois $G$ and $H$, we construct the fiber product $X_{G}\times_{X_{0}(1)}X_{H}$. This is the curve given by $\pi_{G}(x)=\pi_{H}(y)$. The results of this computation are listed in the following table.
@@
Table 1. Potential composite level nilpotent images
The curves that are genus 1 or 2 with rank zero can be handled using standard techniques. The rank 0 genus 3 curve was shown to have no non-cuspidal rational points corresponding to elliptic curves without complex multiplication in [31]. This leaves us with with one remaining curve, with label 35.315.19.1. Theorem A.7 of Appendix A to [33] shows that if $X_{H}$ is a modular curve of level $N$, every simple factor of the Jacobian of $X_{H}$ is isogenous to a simple factor of $J_{1}(N^{2})$, and Section 6 of [33] explains how this information can be used to determine the decomposition of the Jacobian of $X_{H}$. This decomposition is recorded in the beta version of the LMFDB at the link here. In particular, $X_{H}$ factors up to isogeny as the product of nine $\mathbb{Q}$-simple abelian varieties (of dimensions $2$, 33 and $4$).
Work of Kolyvagin-Logachev [21] shows that if a simple abelian variety $A$ of $GL_{2}$-type has analytic rank 0 it must have rank 0, while if it has analytic rank $r$ equal to its dimension, then it must also have algebraic rank $r$. (This latter result is not stated by Kolyvagin and Logachev, but their ideas suffice to prove it. For more detail about how this follows, see [10, Section 7].) This hypothesis is easy to verify since $L(A,s)$ factors as a product of $\dim A$ modular $L$-functions. From this, it follows that the analytic and algebraic rank of the Jacobian of 35.315.19.1 are both $15$. In theory, this curve could be attacked using the method of Chabauty and Coleman since the genus higher than the rank, but computing on a curve of genus 19 is rather unwieldy. However, the techniques of Lemos [25] apply to this situation.
The first theorem proved by Lemos in [25] states the following.
@@
[25, Theorem 1.4] Let $E/\mathbb{Q}$ be an elliptic curve without complex multiplication. Suppose that there exists a prime $q$ for which $\operatorname{Im}\bar{\rho}_{E,q}$ is contained in a subgroup of $C_{s}^{+}(q)$. Then $\bar{\rho}_{E,p}$ is surjective for all $p>37.$
The work in Lemos’s paper does not, however, require full strength of the assumption that $p>37$. We wish to explain why the work in Lemos’s paper implies the following result.
Theorem 6.2.
Let $E/\mathbb{Q}$ be an elliptic curve without complex multiplication. Suppose that there exist a prime $q$ for which $\operatorname{Im}\bar{\rho}_{E,q}$ is contained in a subgroup of $C_{s}^{+}(q)$. Then $\bar{\rho}_{E,p}$ is not contained in $C_{ns}^{+}(p)$ for any $p>3$ with $p\neq q$.
Setting $p=7$ and $q=5$ above implies that the only rational points on the modular curve 35.315.19.1 are cusps or CM points.
Proof of Theorem 6.2.
As this result is really contained in [25] we will give an overview of the steps in Lemos’s argument highlighting the necessary hypotheses on $p$ and $q$. First, Theorem 4.11 implies that $q\in\{2,3,5,7\}$. As a consequence, $X_{0}(q)$ has genus zero.
The modular curve $X_{s}(q)$ parametrizes elliptic curves with two independent cyclic $q$-isogenies. In Lemma 3.4 of [25], Lemos shows that there is an isomorphism
$\theta\colon X_{s}(q)\times_{X_{0}(1)}X_{ns}^{+}(p)\to X_{0}(q^{2})\times_{X_{0}(1)}X_{ns}^{+}(p)$
and that $\theta$ commutes with natural involutions on the source and the target. (The involution on the source interchanges the kernels of the two isogenies, and the involution on the target is the Atkin-Lehner involution $w_{q^{2}}$.) Lemos then defines a map $g\colon X_{0}(q^{2})\times_{X_{0}(1)}X_{ns}^{+}(p)\to J(X_{0}(q)\times_{X_{0}(1)}X_{ns}^{+}(p))$ by taking a point $P$ and mapping it to the difference of its images under the two different degeneracy maps (coming from the two covers $X_{0}(q^{2})\to X_{0}(q)$).
Work of Darmon and Merel [9, Proposition 7.1] shows that there is a projection $\pi\colon J(X_{0}(q)\times_{X_{0}(1)}X_{ns}^{+}(p))\to A$ to a positive-dimensional abelian variety with rank 0 for which the kernel of $\pi$ is connected and is stable under the action of Hecke operators. (For their application, Darmon and Merel only need this for $q=2$ or $q=3$, but as they indicate, this part of the argument applies to any $q$ with $p\nmid q$.)
Define $h=\pi\circ g$ and let $\mathcal{O}$ be the ring of integers in $\mathbb{Q}(\zeta_{p}+\zeta_{p}^{-1})$ and $R=\mathcal{O}[1/(2qp)]$. Roughly speaking, Proposition 3.6 of [25] shows that $h$ is a formal immersion at $\infty$ for every prime ideal $\mathfrak{p}$ of $R$. The proof of this proposition is standard and the key assumption is that $X_{0}(q)$ has genus 0. It follows from this that $\theta\circ h\colon X_{s}(q)\times_{X_{0}(1)}X_{ns}^{+}(p)\to A$ is a formal immersion at $\infty$ for every prime ideal $\mathfrak{p}$ of $R$.
Now, if $P$ is a rational point on $X_{s}(q)\times_{X_{0}(1)}X_{ns}^{+}(p)$ corresponding to an elliptic curve $E$ which has potential multiplicative reduction at some prime $\ell\not\in\{2,p,q\}$, then it meets one of the cusps at the fiber at $\ell$. Without loss of generality this cusp can be chosen to be infinity and this implies that $\theta\circ h(P)=Q\in A(\mathbb{Q})$ reduces to zero in $\tilde{A}(\mathbb{F}_{\ell})$. (Here $\tilde{A}$ is the special fiber of the Néron model of $A$ over $\mathbb{Z}_{\ell}$.) This implies that $Q=0$ and the fact that $\theta\circ h$ is a formal immersion and $\ell>2$ implies that $P=\infty$. (One way to make this conclusion is using Proposition 2.4 of [23].)
It follows from this that if $E/\mathbb{Q}$ is an elliptic curve with mod $p$ image contained in $C_{ns}^{+}(p)$ and mod $q$ image contained in $C_{s}^{+}(q)$, then $E$ cannot have potentially multiplicative reduction at any prime $\ell\not\in\{2,p,q\}$. The primes of potentially multiplicative reduction are precisely those that divide the denominator of $j(E)$. As a consequence $j(E)\in\mathbb{Z}[1/(2pq)]$. More is true however. Lemos notes in [25, Proposition 3.3] that the assumption that the mod $p$ image of Galois is contained in $C_{ns}^{+}(p)$ implies that $E$ has potentially supersingular reduction at $p$, and that if $E$ has potentially multiplicative reduction at $\ell\neq p$, then $\ell\equiv\pm 1\pmod{p}$. (These same observations were made earlier by Zywina.) It follows that the only primes that can divide the denominator of $j(E)$ are those that are $\equiv\pm 1\pmod{p}$. Since $p>3$, none of $2$, 33, 55 or $7$ (the only options for $q$) can be $\equiv\pm 1\pmod{p}$. It follows that $j(E)\in\mathbb{Z}$.
Lemos proceeds to show using an explicit isomorphism $X_{s}^{+}(q)\simeq\mathbb{P}^{1}$ that there are only finite integral $j$-invariants of elliptic curves $E/\mathbb{Q}$ with mod $q$ image of Galois contained in $C_{s}^{+}(q)$. For $q\in\{3,5,7\}$ these are explicitly listed on [25, p. 749]. They all have CM except for j=5000j=-5000 ($q=5$) and j=1728j=-1728 ($q=3$). For elliptic curves with these $j$-invariants, the image of $\rho_{E,p}$ is not contained in the normalizer of a non-split Cartan subgroup for any $p>3$. The case of $q=2$ was previously handled in [24, p. 142]. Here there are $25$ integral $j$-invariants of elliptic curves $E$ with image in $C_{s}^{+}(2)$ (which is a Borel subgroup of $\operatorname{GL}_{2}(\mathbb{F}_{2})$). Of these $25$, there are $18$ are non-CM $j$-invariants. For elliptic curves with these $j$-invariants, the LMFDB indicates that the image of $\rho_{E,p}$ is contained in $C_{ns}^{+}(p)$ for some $p$ only for $p=3$ and $j\in\{-64,4913,238328,16974593\}$. ∎
In Table 2, we give models for the modular curves from Table 1 that have non-cuspidal rational points.
@@ 21.63.1.1 $y^{2}+y=x^{3}+12$ $f_{21}(x,y)=f_{7}\left(\frac{x^{2}+5x-14}{x^{2}-4x+3y+19}\right)$
Table 2. Models of the modular curves relevant for Theorem 1.9. The models for curves of prime level come from [41]. The remaining genus 0 curves can be computed as fiber products of the curves of prime level. The models for the genus 1 modular curves of composite level are computed as the fiber product of the prime level modular curves that are computed in [41] as well.
Working without the assumption of Conjecture 1.8, we must consider the possibility that there is an elliptic curve $E/\mathbb{Q}$ for which $\operatorname{Im}\bar{\rho}_{E,5}$ is contained in the normalizer of the split Cartan mod $5$ and for which $\operatorname{Im}\bar{\rho}_{E,p}$ is contained in the normalizer of a non-split Cartan modulo $p$ for some Mersenne prime $p$. Theorem 6.2 above shows this is only possible for $p=3$, and the elliptic curves for which this occurs are parametrized by the modular curve with label 15.45.1.1.
In addition, we must consider the possibility that there is an elliptic curve $E/\mathbb{Q}$ for which $\mathrm{im}\,\bar{\rho}_{E,p}$ is contained in the normalizer of a non-split Cartan modulo $p$ for some Mersenne prime $p$, and which also has a rational point of order $2$. This implies the mod $2$ image of Galois is contained in a Borel subgroup, which for $p=2$ is equal to $C_{s}^{+}(2)$. The desired result again follows from Theorem 6.2. This concludes the proof of Theorem 1.9.